OpenStudy (anonymous):

Alright I need some help. I really don't know how to do algebra II and I need someone to show me the steps to this problem so hopefully i'll learn. Let f(x) = -5x-4 and g(x) = 6x-7. Find f(x) and g(x).

OpenStudy (diamondboy):

what are u to solve for?

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

Let me take you through the first step, f(x) = -5x -4 g(x) = 6x - 7

OpenStudy (anonymous):

What you do is substitute, here, f(x) = -5(6x - 7) -4 Solve this

OpenStudy (anonymous):

If you still don't understand, let me know

OpenStudy (diamondboy):

did d question say solve for f(g(x))

OpenStudy (anonymous):

f(x) = (-30 + 35) - 4?

OpenStudy (anonymous):

Wait, hold on ._.

OpenStudy (anonymous):

-30x

OpenStudy (anonymous):

This questions seems odd, it says find g(x) and f(x) when it gives us f(x) and g(x). I must be missing something....

OpenStudy (diamondboy):

Yep....d question is wierd

OpenStudy (anonymous):

"Let f(x) = -5x - 4 and g(x) = 6x - 7. Find f(x) + g(x)"

OpenStudy (anonymous):

oooooooooooh

OpenStudy (diamondboy):

Now that is d real question

OpenStudy (anonymous):

Did I type something different??..

OpenStudy (anonymous):

So its (-5x-4) + (6x-7)

OpenStudy (anonymous):

And yeah, you forgot to add the plus sign

OpenStudy (diamondboy):

 yep